# Elimination Distance to Dominated Clusters

### v1

@@ B.1 Marks and unbreakable bag graph
  • for every connected component $C$ of $G-S$ either :
@@
Intuitively, we mark how the set $A$ overlaps with the deleted vertices and the Blue dominating sets. The Blue dominating sets must dominate all Red vertices of $G-A$, and if some Red vertices of $A$ are not dominated, it is written so on $U_{A}$ (for un-dominated).
@@
A $d$-gadget is the graph $H$ composed of $12d+$ 1 vertices $a,b_{1},\ldots,b_{d},b^{\prime}_{1}\ldots,b^{\prime}_{d}$, where for every $i\leq d$, $b_{i}$ is adjacent to $a$ and $b^{\prime}_{i}$.
@@
Let $(T,\mathsf{bag})$ be a regular $(q,k)$-unbreakable tree decomposition of a graph $G$, $S\subseteq\mathsf{cone}(x)$ a set of at most $k$ vertices, and $d$ an integer. We note $\mathsf{bgraph}_{S}(x)$ of a node $x\in V(T)$ any graph that is obtained from $G[\mathsf{cone}(x)-S]$ by :
@@
Every $\mathsf{bgraph}_{S}(x)$ is a $(q,k-|S|,d)$-bag graph.
@@
The vertices of $\mathsf{bgraph}_{S}(x)$ are split into $I$ the vertices of $\mathsf{bag}(x)$, and $E$ the union of all the $d$-gadgets. Clearly, every connected component of $G-I$ is a $d$-gadget and therefore has at most $12d+$ 1 vertices. Furthermore, this $d$-gadget is plugged to a subset of the adhesion of $x$ or the adhesion of a child of $x$, and therefore has at most $q$ neighbors in $I$.
For the last point, let $L,R$ be a separation of $\mathsf{bgraph}_{S}(x)$ of order at most $k-|S|$, then extend $L,R$ into $L^{\prime},R^{\prime}$ by first removing the $d$-gadget plugged to the adhesion of $x$, then adding $S$ to both $L$ and $R$, and finally replacing every $d$-gadget plugged to a subset of $\mathsf{adh}(y)$ by the component of $\mathsf{cone}(y)$ it represents.
@@
For every $x\in V(T)$, $\mathsf{bag}(x)$ is the disjoint union of $\mathsf{adh}(x)$ and $\mathsf{mrg}(x)$. Every vertex of $G$ appears in exactly one margin of one node in the tree decomposition and each node of the tree decomposition has an adhesion of size at most $q$. Furthermore, each node is the child of at most one node and each connected component of a child can be divided in at most $q$ many connected components. Additionally, $d$-gadgets have size at most $12d+$ 1, and $|T|\leq|G|$, hence:
@@
In the next lemma, we show that the ladder index of these bag graphs is bounded (assuming that it was already bounded in the original graph). This together with the fact that the semi-ladder index has 1 implies that the semi-ladder index also has 2.
@@
We state the next proposition with arbitrary parameter $t(\cdot)$ satisfying 2. By the previous lemma, the semi-ladder index $\ell$ can replace $t(\cdot)$.
@@
Given a $(q,k)$-unbreakable tree decomposition of a graph $G$, a node $x$ of the tree decomposition and, for each child $y$ of $x$, the $(k,d)$-profile of $\mathsf{adh}(y)$ in $\mathsf{cone}(y)$, we can compute the $(k,d)$-profile of $\mathsf{adh}(x)$ in $\mathsf{cone}(x)$, in time $f(q,k,d,t(G))\cdot|\mathsf{bgraph}(x)|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying 2. Or alternatively in time $f(q,k,d)\cdot|\mathsf{bgraph}(x)|^{d}$.
Note that by Lemma B.9, the sum of the size of all bag graphs is linear in the size of the graph so that Proposition B.13 yields Theorem 5.4, and therefore Theorem 1.1.
B.2 Algorithm for Proposition B.13
@@
Otherwise, our algorithm branches over all the realized marks of $y$. For each branch, considering a mark $M^{\prime}$, we check the consistency of $M^{\prime}$ with our partial solution, and if it is consistent we update our connected component labeling and our partial solution. The checks are :
@@
  • for every $P\in{\cal{P}}_{x}$ plug the $(d-p)$-gadget to the vertices of $P$.
We are now done building a bag graph $H$, of the form $\mathsf{bgraph}_{S}(x)$. We now conclude by testing whether $H$ is a positive instance to $(k_{x}-|S|-s,d)$-Annotated Dominated Cluster Deletion with Proposition 5.3.
@@
For the first direction, assume that $H$ is a positive instance to $(k_{x}-|S|-s,d)$-Annotated Dominated Cluster Deletion, then the algorithm of Proposition 5.3 produced a set $S^{\prime}$ of size $k_{x}-|S|-s$, which together with $S$ and the sets $S_{y}$ (of cumulative size $s$) selected for the children satisfies that every connected component of $\mathsf{cone}(x)-S-S^{\prime}-\bigcup\limits_{y\in\mathsf{children}(x)}S_{y}$ that does not intersect $\mathsf{adh}(x)$ has a Red-Blue dominating set of size $d$. For the component intersecting $\mathsf{adh}(x)$ the Red-Blue dominating set of size $d$ must contain the $d-p$ Blue vertices of the $(d-p)$-gadget; hence the component has a Red-Blue dominating set of size $p$.
@@ B.3 Running time and Black-White trees
The bound on the number of leafs is straightforward. For the size, notice that whenever a White vertex has some Black children, it only increases the size of the tree to switch them (i.e. the parent of the White vertex now has several Black children, each with one White child), and this operation maintains the degree, budget, and depth of the tree. Therefore, the maximal Black-White tree is a complete Black tree of degree $\alpha$ and of depth $\beta$, where each leaf is attached to a White path of length $n-\beta$, for which the size is at most $\alpha^{\beta}\cdot n$.
When doing the backtracking, it is key that if a child is satisfied with the given partial solution we do not try to branch and make an impactful decision. Otherwise, the tree of partial solutions grows too much and is not linear. In the context of Black-White tree, that would be like allowing one White child and some Black children.
@@
The branching tree of backtracking algorithm of Section B.2 can be labelled to be a Black-White tree of degree $d^{2^{\mathcal{O}(k)}}$ and of budget $(k+qd)$.
@@
When we branch over the children we only have $d^{2^{\mathcal{O}}(k)}$ many marks to consider (see Lemma B.3). This bounds the degree of our Black-White tree.
@@
Going back to the algorithm for Proposition B.13 presented in Section B.2, we just showed that the number of possible branches to consider is only a function of $q,k$, and $d$. And since each branch then perform a single call to Proposition 5.3, this concludes the proof of Theorem 5.4.
@@ Appendix C Annotated version for Elimination Distance to Dominated Clusters
Similarly to Section 5, we extend Theorem 1.2 to an annotated version.
@@
Given a graph $G$, integers $k$, $d$, and vertex sets $F,R,B$, the $(k,d)$-Annotated Elimination Distance to Dominated Clusters problem is to find a tree-structured set $S\subseteq G-F$ of elimination depth $k$ such that every connected component $C$ of $G-S$ has a Red-Blue dominating set of size at most $d$ (i.e. $d$ Blue vertices dominating all Red vertices).
@@
  • for every $i\leq k$ and vertices $u,v$ in $L^{i}$ every $u-v$ path must use a vertex in $L^{j<i}$,
@@
  • for every connected component $C$ of $G-S$ either:
@@
The proof of the next observation is similar to the one of Lemma B.3. {observation} The $(k,d)$-extended-profile of a set $A$ of size at most $q$ in a graph $G$ has size at most $(d+k)^{2^{\mathcal{O}(k)}}$.
We are now going to work with the same definition of gadget and bag graphs (see Definitions B.5 and B.6).
@@
Given a $(q,k)$-unbreakable tree decomposition of a graph $G$, a node $x$ of the decomposition and, for each child $y$ of $x$, the $(k,d)$-extended-profile of $\mathsf{adh}(y)$ in $\mathsf{cone}(y)$, we can compute the $(k,d)$-extended-profile of $\mathsf{adh}(x)$ in $\mathsf{cone}(x)$, in time either $f(q,k,d)\cdot|\mathsf{bgraph}(x)|^{d}$, or $f(q,k,d,t(G))\cdot|\mathsf{bgraph}(x)|^{\mathcal{O}(1)}$ where $t(\cdot)$ is a graph parameter satisfying 2.
Equivalently, the running time is $f(q,k,d,t_{1}(\mathsf{bgraph}(x)))\cdot|\mathsf{bgraph}(x)|^{\mathcal{O}(1)}$ where $t_{1}(\cdot)$ is a graph parameter satisfying 1. Note that by Lemma B.9, the sum of the size of all bag graphs is linear in the size of the graph so Proposition C.4 yields our main result, Theorem 1.2.
C.1 Algorithm for Proposition C.4
@@
${L^{i}_{y}}=\mathsf{adh}(y)\cap S^{i}\leavevmode\nobreak\ \leavevmode\nobreak\ (\text{for all }i\leq k);\leavevmode\nobreak\ \leavevmode\nobreak\ D_{y}=\mathsf{adh}(y)\cap D\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak$
@@
We now have a fixed set $S=\bigcup\limits_{i\leq k}S^{i}$, and similarly to Section B.2, we compute the bag graph $H_{x}=\mathsf{bgraph}_{S}(x)$. Similarly to Appendix A (see Lemma A.3) positive instances have a small Red-Blue dominating set. This transfers to the elimination distance variation, as it was the case in Lemma 4.21 for unbreakable graphs. So the colored version of skeleton-algorithm (see Definition A.7) can also be used here.
@@
This step can be performed in time $f(q,k,d,t_{1}(H_{x}))\cdot|H_{x}|$, where $t_{1}(\cdot)$ satisfies 1, meaning in time $f(q,k,d,t_{2}(G))\cdot|H_{x}|$, where $t_{2}(\cdot)$ satisfies 2. Alternatively, in time $f(q,k,d)\cdot|H_{x}|^{\mathcal{O}(d)}$, in which case we also store the set $D^{\prime}$ that dominates the large connected component. See Lemma 4.25, which can be adapted to bag graphs.
@@
We now compute a partial dominating set in the large connected component. This is achieved as for Corollary 3.4 for which the colors $F,R,B$ are defined as:
@@
First check that in every small connected component $C_{i>0}$, every Red vertex is dominated by $D_{i}$. We then check that the connected components of $\mathsf{cone}(x)$ all have small Red-Blue domination sets. I.e. there could be connected components $C_{i}$, $C_{j}$ of $\mathsf{bag}(x)$ that both have a neighbor in a $d$-gadget in $H_{x}$. We call two such connected components related.
@@
After that, we check the consistency of the $S^{1},\ldots,S^{k}$ i.e. no two vertices of $S^{1}$ are connected in the graphs, then no two vertices of $S^{2}$ can be connected not using $S^{1}$ etc. This can be done in linear time: Simply brute force the paths in the small components, and, in the big component, every two vertices are connected not using $S$ (otherwise the large component is split).
@@ C.2 Correctness and running time
Both are straightforward adaptation of what was done for Annotated Dominated Cluster Deletion. For example Section B.3 also exactly explain the running time of the current algorithm.

### v2

@@ B.1 Marks and unbreakable bag graph
  • for every connected component $C$ of GSG-S either :
@@
Intuitively, we mark how the set $A$ overlaps with the deleted vertices and the Blue dominating sets. The Blue dominating sets must dominate all Red vertices of GAG-A, and if some Red vertices of $A$ are not dominated, it is written so on $U_{A}$ (for undominated).
@@
A $d$-gadget is the graph $H$ composed of $2d+1$ vertices $a,b_{1},\ldots,b_{d},b^{\prime}_{1}\ldots,b^{\prime}_{d}$, where for every $i\leq d$, $b_{i}$ is adjacent to $a$ and $b^{\prime}_{i}$.
@@
Let $(T,\mathsf{bag})$ be a regular $(q,k)$-unbreakable tree decomposition of a graph $G$, $S\subseteq\mathsf{cone}(x)$ a set of at most $k$ vertices, and $d$ an integer. We note $\mathsf{bgraph}_{S}(x)$ of a node $x\in V(T)$ any graph that is obtained from $G[\mathsf{cone}(x)-S]$ by
@@
Every $\mathsf{bgraph}_{S}(x)$ is a (q,k|S|,d)(q,k-|S|,d)-bag graph.
@@
The vertices of $\mathsf{bgraph}_{S}(x)$ are split into $I$ the vertices of $\mathsf{bag}(x)$, and $E$ the union of all the $d$-gadgets. Clearly, every connected component of GIG-I is a $d$-gadget and therefore has at most $2d+1$ vertices. Furthermore, this $d$-gadget is plugged to a subset of the adhesion of $x$ or the adhesion of a child of $x$, and therefore has at most $q$ neighbors in $I$.
For the last point, let $L,R$ be a separation of $\mathsf{bgraph}_{S}(x)$ of order at most k|S|k-|S|, then extend $L,R$ into $L^{\prime},R^{\prime}$ by first removing the $d$-gadget plugged to the adhesion of $x$, then adding $S$ to both $L$ and $R$, and finally replacing every $d$-gadget plugged to a subset of $\mathsf{adh}(y)$ by the component of $\mathsf{cone}(y)$ it represents.
@@
For every $x\in V(T)$, $\mathsf{bag}(x)$ is the disjoint union of $\mathsf{adh}(x)$ and $\mathsf{mrg}(x)$. Every vertex of $G$ appears in exactly one margin of one node in the tree decomposition and each node of the tree decomposition has an adhesion of size at most $q$. Furthermore, each node is the child of at most one node and each connected component of a child can be divided in at most $q$ many connected components. Additionally, $d$-gadgets have size at most $2d+1$, and $|T|\leq|G|$, hence:
@@
In the next lemma, we show that the ladder index of these bag graphs is bounded (assuming that it was already bounded in the original graph). This together with the fact that the semi-ladder index has ˜1 implies that the semi-ladder index also has ˜2.
@@
We state the next proposition with arbitrary parameter $t(\cdot)$ satisfying ˜2. By the previous lemma, the semi-ladder index $\ell$ can replace $t(\cdot)$.
@@
Given a $(q,k)$-unbreakable tree decomposition of a graph $G$, a node $x$ of the tree decomposition and, for each child $y$ of $x$, the $(k,d)$-profile of $\mathsf{adh}(y)$ in $\mathsf{cone}(y)$, we can compute the $(k,d)$-profile of $\mathsf{adh}(x)$ in $\mathsf{cone}(x)$, in time $f(q,k,d,t(G))\cdot|\mathsf{bgraph}(x)|^{\mathcal{O}(1)}$, where $t(\cdot)$ is a parameter satisfying ˜2, or alternatively in time $f(q,k,d)\cdot|\mathsf{bgraph}(x)|^{d}$.
Note that by Lemma˜B.9, the sum of the size of all bag graphs is linear in the size of the graph so that Proposition˜B.13 yields Theorem˜5.4, and therefore Theorem˜1.1.
B.2 Algorithm for Proposition˜B.13
@@
Otherwise, our algorithm branches over all the realized marks of $y$. For each branch, considering a mark $M^{\prime}$, we check the consistency of $M^{\prime}$ with our partial solution, and if it is consistent, we update our connected component labeling and our partial solution.
The checks are:
@@
  • for every $P\in{\cal{P}}_{x}$ plug the (dp)(d-p)-gadget to the vertices of $P$.
We are now done building a bag graph $H$, of the form $\mathsf{bgraph}_{S}(x)$. We now conclude by testing whether $H$ is a positive instance to $(k_{x}-|S|-s,d)$-Annotated Dominated Cluster Deletion with Proposition˜5.3.
@@
For the first direction, assume that $H$ is a positive instance to $(k_{x}-|S|-s,d)$-Annotated Dominated Cluster Deletion, then the algorithm of Proposition˜5.3 produced a set $S^{\prime}$ of size $k_{x}-|S|-s$, which together with $S$ and the sets $S_{y}$ (of cumulative size $s$) selected for the children satisfies that every connected component of $\mathsf{cone}(x)-S-S^{\prime}-\bigcup\limits_{y\in\mathsf{children}(x)}S_{y}$ that does not intersect $\mathsf{adh}(x)$ has a Red-Blue dominating set of size $d$. For the component intersecting $\mathsf{adh}(x)$ the Red-Blue dominating set of size $d$ must contain the dpd-p Blue vertices of the (dp)(d-p)-gadget; hence the component has a Red-Blue dominating set of size $p$.
@@ B.3 Running time and Black-White trees
The bound on the number of leaves is straightforward. For the size, notice that whenever a White vertex has some Black children, it only increases the size of the tree to switch them (i.e. the parent of the White vertex now has several Black children, each with one White child), and this operation maintains the degree, budget, and depth of the tree. Therefore, the maximal Black-White tree is a complete Black tree of degree $\alpha$ and of depth $\beta$, where each leaf is attached to a White path of length $n-\beta$, for which the size is at most $\alpha^{\beta}\cdot n$.
When doing the backtracking, it is key that if a child is satisfied with the given partial solution, we do not try to branch and make an impactful decision. Otherwise, the tree of partial solutions grows too much and is not linear. In the context of Black-White tree, that would be like allowing one White child and some Black children.
@@
The branching tree of the backtracking algorithm of Section˜B.2 can be labelled to be a Black-White tree of degree $d^{2^{\mathcal{O}(k)}}$ and of budget $(k+qd)$.
@@
When we branch over the children we only have $d^{2^{\mathcal{O}}(k)}$ many marks to consider (see Lemma˜B.3). This bounds the degree of our Black-White tree.
@@
Going back to the algorithm for Proposition˜B.13 presented in Section˜B.2, we just showed that the number of possible branches to consider is only a function of $q,k$, and $d$. And since each branch then perform a single call to Proposition˜5.3, this concludes the proof of Theorem˜5.4.
@@ Appendix C Annotated version for Elimination Distance to Dominated Clusters
Similarly to Section˜5, we extend Theorem˜1.2 to an annotated version.
@@
Given a graph $G$, integers $k$, $d$, and vertex sets $F,R,B$, the $(k,d)$-Annotated Elimination Distance to Dominated Clusters problem is to find a tree-structured set $S\subseteq G-F$ of elimination depth $k$ such that every connected component $C$ of GSG-S has a Red-Blue dominating set of size at most $d$ (i.e. $d$ Blue vertices dominating all Red vertices).
@@
  • for every $i\leq k$ and vertices $u,v$ in $L^{i}$ every uvu-v path must use a vertex in $L^{j<i}$,
@@
  • for every connected component $C$ of GSG-S either:
@@
The proof of the next observation is similar to the one of Lemma˜B.3. {observation} The $(k,d)$-extended-profile of a set $A$ of size at most $q$ in a graph $G$ has size at most $(d+k)^{2^{\mathcal{O}(k)}}$.
We are now going to work with the same definition of gadget and bag graphs (see Definitions˜B.5 and B.6).
@@
Given a $(q,k)$-unbreakable tree decomposition of a graph $G$, a node $x$ of the decomposition and, for each child $y$ of $x$, the $(k,d)$-extended-profile of $\mathsf{adh}(y)$ in $\mathsf{cone}(y)$, we can compute the $(k,d)$-extended-profile of $\mathsf{adh}(x)$ in $\mathsf{cone}(x)$, in time either $f(q,k,d)\cdot|\mathsf{bgraph}(x)|^{d}$, or $f(q,k,d,t(G))\cdot|\mathsf{bgraph}(x)|^{\mathcal{O}(1)}$ where $t(\cdot)$ is a graph parameter satisfying ˜2.
Equivalently, the running time is $f(q,k,d,t_{1}(\mathsf{bgraph}(x)))\cdot|\mathsf{bgraph}(x)|^{\mathcal{O}(1)}$ where $t_{1}(\cdot)$ is a graph parameter satisfying ˜1. Note that by Lemma˜B.9, the sum of the size of all bag graphs is linear in the size of the graph so Proposition˜C.4 yields our main result, Theorem˜1.2.
C.1 Algorithm for Proposition˜C.4
@@
${L^{i}_{y}}=\mathsf{adh}(y)\cap S^{i}\leavevmode\nobreak\ \leavevmode\nobreak\ (\text{for all }i\leq k);\leavevmode\nobreak\ \leavevmode\nobreak\ D_{y}=\mathsf{adh}(y)\cap D\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\$
@@
We now have a fixed set $S=\bigcup\limits_{i\leq k}S^{i}$, and similarly to Section˜B.2, we compute the bag graph $H_{x}=\mathsf{bgraph}_{S}(x)$. Similarly to Appendix˜A (see Lemma˜A.3) positive instances have a small Red-Blue dominating set. This transfers to the elimination distance variation, as it was the case in Lemma˜4.21 for unbreakable graphs. So the colored version of skeleton-algorithm (see Definition˜A.7) can also be used here.
@@
This step can be performed in time $f(q,k,d,t_{1}(H_{x}))\cdot|H_{x}|$, where $t_{1}(\cdot)$ satisfies ˜1, meaning in time $f(q,k,d,t_{2}(G))\cdot|H_{x}|$, where $t_{2}(\cdot)$ satisfies ˜2. Alternatively, in time $f(q,k,d)\cdot|H_{x}|^{\mathcal{O}(d)}$, in which case we also store the set $D^{\prime}$ that dominates the large connected component. See Lemma˜4.25, which can be adapted to bag graphs.
@@
We now compute a partial dominating set in the large connected component. This is achieved as for Corollary˜3.4 for which the colors $F,R,B$ are defined as:
@@
First check that in every small connected component $C_{i>0}$, every Red vertex is dominated by $D_{i}$. We then check that the connected components of $\mathsf{cone}(x)$ all have small Red-Blue domination sets, i.e. there could be connected components $C_{i}$, $C_{j}$ of $\mathsf{bag}(x)$ that both have a neighbor in a $d$-gadget in $H_{x}$. We call two such connected components related.
@@
After that, we check the consistency of the $S^{1},\ldots,S^{k}$, i.e. no two vertices of $S^{1}$ are connected in the graphs, then no two vertices of $S^{2}$ can be connected not using $S^{1}$ etc. This can be done in linear time: Simply brute force the paths in the small components, and, in the big component, every two vertices are connected not using $S$ (otherwise the large component is split).
@@ C.2 Correctness and running time
Both are straightforward adaptation of what was done for Annotated Dominated Cluster Deletion. For example, Section˜B.3 also exactly explains the running time of the current algorithm.